A body is thrown with an initial velocity v0 at an angle to the horizon . The duration of the motion in 2.2 seconds . The starting point A is on horiontal ground and B is the point where it lands on the ground . Find the maximum height reached by the body
 

Grade:11

1 Answers

Aman gohel
6 years ago
angle=X,u=V0,T=2.2.Time=(2usinX)/g=>2.2=(2*V0*sinX)/9.8=>V0sinX=10.78=>V0^2*sinX^2=116.20Height=u^2sin^2X/2g=>116.20/19.8Height(max)=5.929m(Ans)